# TESS discovery of two super-Earths orbiting the M-dwarf stars TOI-6002 and TOI-5713 near the radius valley

### v1

@@
TESS is increasing the number of small exoplanets amenable to mass and atmospheric characterization (Ricker et al., 2015; Sullivan et al., 2015). Detailed studies of such exoplanets will allow insights into the origin of the radius/density valley in the near future. Such analyses should shed further light on the mechanisms dominating the formation and evolution of super-Earths. In addition the analysis of exoplanets with incident irradiation between Earth and Venus will allow key insights into the limits of the HZ (e.g. Kaltenegger et al., 2023a).
@@ 2 Observations
. Date (UT) Filter Facility Pixel scale (″/pixel) Exp. time [s] N. of exposures Airmass range Aperture size (″) Detrending parameters TOI-6002 13 May 2023 LCO-SINISTRO (1.0 m) 0.389 86 93 [1.00,1.40] 2.33 Airmass+Background 24 May 2023 LCO-SINISTRO (1.0 m) 0.389 86 83 [1.04,1.89] 2.33 Airmass+FWHM 19 August 2023 LCOGT MuSCAT3 (2.0 m) 0.265 46, 32 251, 339 [1.03,1.30] 2.38 Airmass+FWHM – 207, 28 35, 375 – – Airmass+Background TOI-5713 17 June 2023 TRAPPIST-North (0.6 m) 0.640 50 209 [1.25,2.88] 5.40 Airmass+FWHM 08 July 2023 TCS-MuSCAT2 (1.52 m) 0.440 50, 25 195, 457 [1.16,1.94] 6.96 Airmass+FWHM – 15 234 – – Airmass+Background
@@ 2.5 Gaia Assessment
We also used the Gaia renormalised unit weight error (RUWE) to assess the possibility of any close-in companions. RUWE values indicate excess astrometric noise, which can potentially be caused by a bound companion (Ziegler et al., 2020). The Gaia DR3 RUWE values for TOI-5713 and TOI-6002 are 1.3 and 1.1, respectively, indicating astrometric solutions that are consistent with single-star models.
@@ 3.2 Discarding possible false-positive scenarios
Two primary cases of astrophysical false positives can masquerade as exoplanet transits in TESS data. The first case is background/foreground eclipsing binaries (EBs), especially those in grazing configurations. For both targets, we searched for archival imaging, retrieving images from POSS I/DSS (Minkowski & Abell, 1963) and POSS II/DSS Lasker et al. (1996). These images, in addition to those taken from our recent observations from MusCAT3 (for TOI-6002) and MusCAT2 (for TOI-5713) in 2023, span more than 70 years, as shown in Figure 7. TOI-6002 has a high proper motion of 344.42 mas/yr and has moved by 24.4″ from 1955 to 2023. This allowed us to confirm that there was no bright star in the current position of TOI-6002. On the other hand, TOI-5713 has a relatively low proper motion (110.24 mas/yr) and has moved by 7.7″ from 1953 to 2023, making it challenging to confirm the absence of any unresolved companion in its current position. However, high-resolution adaptive optics imaging from Keck-II (for TOI-5713) and Palomar (for TOI-6002) confirm that no stellar companions were detected close to () the targets as discussed in Sect. 2.4. Additionally, the renormalized unit weight error (RUWE) from Gaia DR3 is below 1.4 for both stars, indicating that the stars are singles (see Sect. 2.5).
@@ 3.4 Planet search and detection limits
As mentioned in Section 2.1, TOI-5713 and TOI-6002 have both been observed in multiple sectors of TESS — respectively, sectors 16, 22, 23, 46, and sectors 14, 41, 54, 55. Since both planets were detected by the TESS pipelines with a low signal-to-noise ratio, it is possible additional planets remained unnoticed below the threshold set to trigger an alert (Multiple Event Statistic = 7.1) (Tenenbaum et al., 2012). In this context, we performed a search for additional planet candidates in the TESS data with the SHERLOCK666https://github.com/franpoz/SHERLOCK pipeline (Pozuelos et al., 2020; Demory et al., 2020; Dévora-Pajares et al., 2024) and assessed the detection limits with injection and recovery tests with the MATRIX777https://github.com/PlanetHunters/tkmatrix ToolKit (Pozuelos et al., 2020; Dévora-Pajares & Pozuelos, 2022).
@@
The lack of additional transiting planet candidates in the TESS data for both systems does not necessarily exclude the existence of extra planets. Several scenarios could explain the lack of interesting signals in our transit search with SHERLOCK: (1) the systems only host one planet; (2) if other planets exist, they are not transiting; (3) the data we examined do not have the photometric precision required to detect the additional transiting planets, or they have a longer period than the period grid searched in our analyses (Wells et al., 2021). In the case of scenario (2), radial velocity measurements may be able to detect them, as discussed in Section 4.2. In the case of scenario (3), a ground-based high-precision photometric follow-up could possibly detect more planets, such as was the case for LP 890-9 c (Delrez et al., 2022). However, to test scenario (3) against the TESS data, we performed injection and recovery tests over the available short-cadence data for both systems.
@@
Our tests show that TOI-5713 b and TOI-6002 b lie at the detectability limits found for each of their systems (see Figure 8). This is consistent with the low S/N obtained for both planets, either in SHERLOCK (, ) or SPOC (, ), and the low number of SHERLOCK detrends where they were found. It is essential to recognize that the S/N reported by the SHERLOCK algorithm is derived from the TLS package, which employs a straightforward estimation approach that is not directly comparable to the metrics produced by alternative pipelines, such as SPOC. Instead, SHERLOCK utilizes these values internally to evaluate the detected signals and identify the most significant ones among them.
@@ 3.5 Stellar activity and rotation period
We then used the Systematics-Insensitive Periodogram (TESS-SIP888https://github.com/christinahedges/TESS-SIP) (Hedges et al., 2020) to build a periodogram using the whole TESS photometry (4 sectors) for each of the two targets. This tool simultaneously creates a Lomb-Scargle periodogram and detrends TESS systematics as described in Angus et al. (2016). The SPOC pipeline is known to attenuate rotational periods, thus TESS-SIP reproduces simple aperture photometry (SAP) light curves from the TPFs and the SPOC apertures. We applied this procedure on all four sectors observed by TESS. Searching for periods between 2 and 15 days, we applied TESS-SIP on the target and on all the background pixels around the SPOC apertures. We found no significant rotational signal for TOI-6002; however, we found a significant rotational period at 9.320.20 days for TOI-5713. Thus, we report the SIP power spectrum for only TOI-5713 in Figure 9. We also found that this signal remains on the PDC-SAP light curves of each individual sector, which underscores the robustness of the signal. Beyond 15 days, we did not find any rotational signal.
@@ 4.1 Radius valley
Observational and theoretical studies showed that the position of the radius valley is period-dependent (see Figure 12). Its location has been found trending to smaller planet radii for long periods in both thermally-driven mass loss and gas-poor formation models for solar FGK stars (see e.g. Fulton et al., 2017; Van Eylen et al., 2018; Lopez & Rice, 2018; Lee & Connors, 2021). For M dwarfs, its location has been found trending to larger radii for long periods in the gas-depleted theory (see e.g Lopez & Rice, 2018; Cloutier & Menou, 2020). Conversely, other studies on small planets around M dwarfs found a negative slope (see e.g. Van Eylen et al., 2021; Bonfanti et al., 2023), something similar to that of FGK stars.
As shown in Figure 12, a wedge of opposing predictions emerged because of the difference in the slopes of the radius valley. Planets in this region are known as ”keystone planets”. They are predicted to be purely rocky by thermally-driven mass loss and gas-poor formation, and potentially rocky with small gaseous envelopes by the gas-depleted formation model (see e.g. Cherubim et al., 2023). Only a few planets with measured densities are in this wedge. Having a large sample of keystone planets is important to understand the models dominating the formation and evolution of small planets. TOI-5713 b ( and d) is located within the wedge of keystone planets with the possibility of being at its upper edge, given the uncertainty on the planetary radius. TOI-6002 b ( and d) is located at the lower edge of this wedge with the possibility of being inside it, given the uncertainty on the planetary radius. When their masses are determined with RV observations (see Sect.4.2), these planets will join the small sample of keystone planets with density measurements. This sample will, in the future, help disentangle which model is dominating the formation of small (i.e., ) planets.
@@ 4.3 Potential for atmospheric characterization
Ehrenreich et al. (2012) showed that during Venus’s 2012 transit, its transmission spectra indicated that the cloud and haze layers obstruct probing below an altitude of 80 km. Pidhorodetska et al. (2021) modeled the transmission spectra of the L 98-59 planets, a benchmark system for studying the VZ planets (Ostberg et al., 2023). They showed that if the L 98-59 planets possess a clear Venus-like atmosphere, NIRSpec could detect CO2 within 26 transits for each planet. However, the presence of H2SO4 clouds would significantly suppress CO2 absorption. A new study by Kaltenegger et al. (2023a) shows that there is a wide range of possible atmospheres for a rocky planet experiencing a runaway greenhouse effect, which changes the atmosphere and observable features significantly. Model transmission spectra of the planet LP 890-9c (Delrez et al., 2022), another benchmark system for studying the VZ planets (Kaltenegger et al., 2023a) show that depending on the state of the planet, JWST observations could infer evidence of H2O with 3 transit (at 3 confidence) for a full runaway greenhouse scenario for LP 890-9c, and CO2-dominated atmospheres resembling Venus without high-altitude terminator clouds with 8 transits (Gomez Barrientos et al., 2023). A 3D Venus model for the same planet with higher, thick cloud coverage increases the amount of transits needed to be able to explore the atmospheric composition. Note that all predictions could be complicated by the impact of clouds and/or unocculted starspots, especially for cool host stars.
@@
These models show that TOI-6002b and TOI-5713b represent promising targets for atmospheric characterization with JWST, as they may provide valuable insights into the diverse evolutionary pathways and climate states of hot rocky or water-rich exoplanets. Note that we assume rocky or water-rich compositions in our models based on the radius of the planets. Mass measurements could help provide better constraints on the viability of atmospheric studies and would be possible with the MAROON-X instrument at the 8.1-m Gemini North telescope.
@@ Acknowledgements.
This research was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). This article is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide. TRAPPIST-North is a project funded by the University of Liege (Belgium), in collaboration with Cadi Ayyad University of Marrakesh (Morocco). We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. KAC acknowledges support from the TESS mission via subaward s3449 from MIT. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This paper is based on observations made with the MuSCAT3 instrument, developed by the Astrobiology Center and under financial supports by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), at Faulkes Telescope North on Maui, HI, operated by the Las Cumbres Observatory. This work was partially supported by a grant from the Erasmus+ International Credit Mobility program (M. Ghachoui). This publication benefits from the support of the French Community of Belgium in the context of the FRIA Doctoral Grant awarded to MT. MG is F.R.S-FNRS Research Director. The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. Author F.J.P acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and Ministerio de Ciencia e Innovación through the project PID2022-137241NB-C43. H.P. acknowledges support by the Spanish Ministry of Science and Innovation with the Ramon y Cajal fellowship number RYC2021-031798-I. This work is partly supported by JSPS KAKENHI Grant Number JP18H05439 and JST CREST Grant Number JPMJCR1761. E.E-B. acknowledges financial support from the European Union and the State Agency of Investigation of the Spanish Ministry of Science and Innovation (MICINN) under the grant PRE2020-093107 of the Pre-Doc Program for the Training of Doctors (FPI-SO) through FSE funds We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This material is based upon work supported by the National Aeronautics and Space Administration under Agreement No. 80NSSC21K0593 for the program “Alien Earths”. The results reported herein benefitted from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate. This research was supported by Wallonia-Brussels InternationalReferences

### v2

@@
TESS is increasing the number of small exoplanets amenable to mass and atmospheric characterization (Ricker et al., 2015; Sullivan et al., 2015). Detailed studies of such exoplanets will allow insights into the origin of the radius/density valley in the near future. Such analyses should shed further light on the mechanisms dominating the formation and evolution of super-Earths. In addition, the analyses of exoplanets with incident irradiation between Earth and Venus will allow key insights into the limits of the HZ (e.g. Kaltenegger et al., 2023a).
@@ 2 Observations
. Date (UT) Filter Facility Pixel Scale (″/pixel) Exposure Time [s] Number of Exposures Airmass Range Aperture Size (″) Detrending Parameters TOI-6002 13 May 2023 LCO-SINISTRO (1.0 m) 0.389 86 93 [1.00,1.40] 2.33 Airmass+Background 24 May 2023 LCO-SINISTRO (1.0 m) 0.389 86 83 [1.04,1.89] 2.33 Airmass+FWHM 19 August 2023 LCOGT MuSCAT3 (2.0 m) 0.265 46, 32 251, 339 [1.03,1.30] 2.38 Airmass+FWHM … 207, 28 35, 375 … … Airmass+Background TOI-5713 17 June 2023 TRAPPIST-North (0.6 m) 0.640 50 209 [1.25,2.88] 5.40 Airmass+FWHM 08 July 2023 TCS-MuSCAT2 (1.52 m) 0.440 50, 25 195, 457 [1.16,1.94] 6.96 Airmass+FWHM … 15 234 … … Airmass+Background
@@ 2.5 Gaia Assessment
We also used the Gaia renormalized unit weight error (RUWE) to assess the possibility of any close-in companions. RUWE values indicate excess astrometric noise, which can potentially be caused by a bound companion (Ziegler et al., 2020). The Gaia DR3 RUWE values for TOI-5713 and TOI-6002 are 1.3 and 1.1, respectively, indicating astrometric solutions that are consistent with single-star models.
@@ 3.2 Discarding possible false-positive scenarios
Two primary cases of astrophysical false positives can masquerade as exoplanet transits in TESS data. The first case is background/foreground eclipsing binaries (EBs), especially those in grazing configurations. For both targets, we searched for archival imaging, retrieving images from POSS I/DSS (Minkowski & Abell, 1963) and POSS II/DSS Lasker et al. (1996). These images, in addition to those taken from our recent observations from MusCAT3 (for TOI-6002) and MusCAT2 (for TOI-5713) in 2023, span more than 70 years, as shown in Fig. 7. TOI-6002 has a high proper motion of 344.42 mas/yr and has moved by 24.4″ from 1955 to 2023. This allowed us to confirm that there was no bright star in the current position of TOI-6002. On the other hand, TOI-5713 has a relatively low proper motion (110.24 mas/yr) and has moved by 7.7″ from 1953 to 2023, making it challenging to confirm the absence of any unresolved companion in its current position. However, high-resolution adaptive optics imaging from Keck-II (for TOI-5713) and Palomar (for TOI-6002) confirm that no stellar companions were detected close to () the targets as discussed in Sect. 2.4. Additionally, the renormalized unit weight error (RUWE) from Gaia DR3 is below 1.4 for both stars, indicating that the stars are singles (see Sect. 2.5).
@@ 3.4 Planet search and detection limits
As mentioned in Sect. 2.1, TOI-5713 and TOI-6002 have both been observed in multiple sectors of TESS — respectively, sectors 16, 22, 23, 46, and sectors 14, 41, 54, 55. Since both planets were detected by the TESS pipelines with a low signal-to-noise ratio, it is possible additional planets remained unnoticed below the threshold set to trigger an alert (Multiple Event Statistic = 7.1) (Tenenbaum et al., 2012). In this context, we performed a search for additional planet candidates in the TESS data with the SHERLOCK666https://github.com/franpoz/SHERLOCK pipeline (Pozuelos et al., 2020; Demory et al., 2020; Dévora-Pajares et al., 2024) and assessed the detection limits with injection and recovery tests with the MATRIX777https://github.com/PlanetHunters/tkmatrix ToolKit (Pozuelos et al., 2020; Dévora-Pajares & Pozuelos, 2022).
@@
The lack of additional transiting planet candidates in the TESS data for both systems does not necessarily exclude the existence of extra planets. Several scenarios could explain the lack of interesting signals in our transit search with SHERLOCK: (1) the systems only host one planet; (2) if other planets exist, they are not transiting; (3) the data we examined do not have the photometric precision required to detect the additional transiting planets, or they have a longer period than the period grid searched in our analyses (Wells et al., 2021). In the case of scenario (2), radial velocity measurements may be able to detect them, as discussed in Sect. 4.2. In the case of scenario (3), a ground-based high-precision photometric follow-up could possibly detect more planets, such as was the case for LP 890-9 c (Delrez et al., 2022). However, to test scenario (3) against the TESS data, we performed injection and recovery tests over the available short-cadence data for both systems.
@@
Our tests show that TOI-5713 b and TOI-6002 b lie at the detectability limits found for each of their systems (see Fig. 8). This is consistent with the low S/N obtained for both planets, either in SHERLOCK (, ) or SPOC (, ), and the low number of SHERLOCK detrends where they were found. It is essential to recognize that the S/N reported by the SHERLOCK algorithm is derived from the TLS package, which employs a straightforward estimation approach that is not directly comparable to the metrics produced by alternative pipelines, such as SPOC. Instead, SHERLOCK utilizes these values internally to evaluate the detected signals and identify the most significant ones among them.
@@ 3.5 Stellar activity and rotation period
We then used the Systematics-Insensitive Periodogram (TESS-SIP888https://github.com/christinahedges/TESS-SIP) (Hedges et al., 2020) to build a periodogram using the whole TESS photometry (4 sectors) for each of the two targets. This tool simultaneously creates a Lomb-Scargle periodogram and detrends TESS systematics as described in Angus et al. (2016). The SPOC pipeline is known to attenuate rotational periods, thus TESS-SIP reproduces simple aperture photometry (SAP) light curves from the TPFs and the SPOC apertures. We applied this procedure on all four sectors observed by TESS. Searching for periods between 2 and 15 days, we applied TESS-SIP on the target and on all the background pixels around the SPOC apertures. We found no significant rotational signal for TOI-6002; however, we found a significant rotational period at 9.320.20 days for TOI-5713. Thus, we report the SIP power spectrum for only TOI-5713 in Fig. 9. We also found that this signal remains on the PDC-SAP light curves of each individual sector, which underscores the robustness of the signal. Beyond 15 days, we did not find any rotational signal.
@@ 4.1 Radius valley
Observational and theoretical studies showed that the position of the radius valley is period-dependent (see Fig. 12). Its location has been found trending to smaller planet radii for long periods in both thermally-driven mass loss and gas-poor formation models for solar FGK stars (see e.g. Fulton et al., 2017; Van Eylen et al., 2018; Lopez & Rice, 2018; Lee & Connors, 2021). For M dwarfs, its location has been found trending to larger radii for long periods in the gas-depleted theory (see e.g Lopez & Rice, 2018; Cloutier & Menou, 2020). Conversely, other studies on small planets around M dwarfs found a negative slope (see e.g. Van Eylen et al., 2021; Bonfanti et al., 2023), something similar to that of FGK stars.
As shown in Fig. 12, a wedge of opposing predictions emerged because of the difference in the slopes of the radius valley. Planets in this region are known as ”keystone planets”. They are predicted to be purely rocky by thermally-driven mass loss and gas-poor formation, and potentially rocky with small gaseous envelopes by the gas-depleted formation model (see e.g. Cherubim et al., 2023). Only a few planets with measured densities are in this wedge. Having a large sample of keystone planets is important to understand the models dominating the formation and evolution of small planets. TOI-5713 b ( and d) is located within the wedge of keystone planets with the possibility of being at its upper edge, given the uncertainty on the planetary radius. TOI-6002 b ( and d) is located at the lower edge of this wedge with the possibility of being inside it, given the uncertainty on the planetary radius. When their masses are determined with RV observations (see Sect.4.2), these planets will join the small sample of keystone planets with density measurements. This sample will, in the future, help disentangle which model is dominating the formation of small (i.e., ) planets.
@@ 4.3 Potential for atmospheric characterization
Ehrenreich et al. (2012) showed that during Venus’s 2012 transit, its transmission spectra indicated that the cloud and haze layers obstruct probing below an altitude of 80 km. Pidhorodetska et al. (2021) modeled the transmission spectra of the L 98-59 planets, a benchmark system for studying the VZ planets (Ostberg et al., 2023). They showed that if the L 98-59 planets possess a clear Venus-like atmosphere, NIRSpec could detect CO2 within 26 transits for each planet. However, the presence of H2SO4 clouds would significantly suppress CO2 absorption. A new study by Kaltenegger et al. (2023a) shows that there is a wide range of possible atmospheres for a rocky planet experiencing a runaway greenhouse effect, which changes the atmosphere and observable features significantly. Model transmission spectra of the planet LP 890-9c (Delrez et al., 2022), another benchmark system for studying the VZ planets (Kaltenegger et al., 2023a) show that depending on the state of the planet, James Webb Space Telescope (JWST) observations could infer evidence of H2O with 3 transit (at 3 confidence) for a full runaway greenhouse scenario for LP 890-9c, and CO2-dominated atmospheres resembling Venus without high-altitude terminator clouds with 8 transits (Gomez Barrientos et al., 2023). A 3D Venus model for the same planet with higher, thick cloud coverage increases the amount of transits needed to be able to explore the atmospheric composition. Note that all predictions could be complicated by the impact of clouds and/or unocculted starspots, especially for cool host stars.
@@
These models show that TOI-6002b and TOI-5713b represent promising targets for atmospheric characterization with JWST, as they may provide valuable insights into the diverse evolutionary pathways and climate states of hot rocky or water-rich exoplanets. Mass measurements could help provide better constraints on the viability of atmospheric studies and would be possible with the MAROON-X instrument at the 8.1-m Gemini North telescope.
@@ Acknowledgements.
This research was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). This article is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide. TRAPPIST-North is a project funded by the University of Liege (Belgium), in collaboration with Cadi Ayyad University of Marrakesh (Morocco). We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. KAC acknowledges support from the TESS mission via subaward s3449 from MIT. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This paper is based on observations made with the MuSCAT3 instrument, developed by the Astrobiology Center and under financial supports by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), at Faulkes Telescope North on Maui, HI, operated by the Las Cumbres Observatory. This work was partially supported by a grant from the Erasmus+ International Credit Mobility program (M. Ghachoui). This publication benefits from the support of the French Community of Belgium in the context of the FRIA Doctoral Grant awarded to MT. MG is F.R.S-FNRS Research Director. The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. Author F.J.P acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and Ministerio de Ciencia e Innovación through the project PID2022-137241NB-C43. H.P. acknowledges support by the Spanish Ministry of Science and Innovation with the Ramon y Cajal fellowship number RYC2021-031798-I. This work is partly supported by JSPS KAKENHI Grant Number JP18H05439 and JST CREST Grant Number JPMJCR1761. E.E-B. acknowledges financial support from the European Union and the State Agency of Investigation of the Spanish Ministry of Science and Innovation (MICINN) under the grant PRE2020-093107 of the Pre-Doc Program for the Training of Doctors (FPI-SO) through FSE funds We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This material is based upon work supported by the National Aeronautics and Space Administration under Agreement No. 80NSSC21K0593 for the program “Alien Earths”. The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate. This research was supported by Wallonia-Brussels International.References